# A new foundation of quantum decision theory

### v4

@@ Abstract
Quantum decision theory is introduced here, and new basis for this theory is proposed. It is first based upon the author’s general arguments for the Hilbert space formalism in quantum theory, next on arguments for the Born rule, that is, the basis for calculating quantum probabilities. A basic notion behind the quantum theory foundation is that of theoretical variables, that are divided into accessible and inaccessible ones. This is here specialized to decision variables. It is assumed that all accessible variables can be seen as functions of a specific inaccessible variable. Another assumption is that there exist two maximal accessible theoretical variables in the given situation. Two basic assumptions behind the Born rule are 1) the likelihood principle, 2) the actor in question has motivations that can be modeled by a hypothetical perfectly rational higher being. The theory is illustrated by a medical example. Finally, a broad discussion of decision processes is given.
@@ 1 Introduction
In standard economic theory and statistical theory, the problem of optimizing behaviour is reduced to optimizing expected reward or utility. I will claim that this does not quite correspond to the way people act during a decision process. This statement can be supported by a number of empirical investigations; see for instance the review articles mentioned below in connection to quantum decision theory. The argument will be further elucidated later in this article.
@@
A rather primitive form of learning is reinforcement learning, learning governed by rewards or punishments. In Friston et al. (2009) the notion of active inference was introduced as an alternative to reinforcement learning in connection to decision making. This is important, since the way people learn in connection to decisions obviously is more active than what can be described as reinforcement learning. In op cit. a theory of active inference is formalized in terms of a free-energy formulation of perception, quantizised in terms of entropy or average surprise associated with a probability distribution on an agent’s state space and his environment.
@@
Theories of decisions can take many points of departure. In this article we will describe a new foundation of a comparatively recent theory: Quantum decision theory (QDT). This theory is completely based upon quantum mechanics, whose foundation is discussed in the book Helland (2021) and in some newer articles, Helland (2022a,b,c, 2023a,b,c, 2024a,b,c).
@@
The plan of this article is as follows: I first review briefly a new approach to the foundation of quantum theory, and then I use this to derive and motivate a new foundation of quantum decision theory, taking as a basis a simple model of a person’s decisions. First, the Hilbert space formalism of decision variables is derived. Then the Born formula, the basis for computing quantum probabilities, is argued for under weak assumptions, and this is specialized to a decision setting. An example of applications is then given, and used to illustrate some of the peculiarities of quantum probabilities. Then my foundation of QDT is further discussed from several points of view before I give some concluding remarks.
@@ 2.1 On the foundation of quantum theory in general
A completely new approach towards quantum foundations is proposed in Helland (2021, 2022a,b,c 2023a,b,c,2024a,b,c). The basis of this approach is an observer who is in some physical situation. In this situation there are theoretical variables, and certain of these variables, say $\theta,\lambda,\eta,...$ are related to the observer $C$. Some of the variables are accessible to him, meaning roughly that it is, in some future, in principle possible to obtain as accurate values as he wishes on the relevant variable. Others are inaccessible. Examples of the latter are the vector (position, momentum) of a particle at some time, or the full spin vector of a spin particle, an imagined vector whose discretized projection in the direction $a$ is the spin component in that direction. The terms ‘accessible’ and ‘inaccessible’ are primitive notions of the theory.
The main assumption of my theory is then as follows: Related to an observer $C$ there is an inaccessible variable $\phi$ such that all accessible variables can be seen as functions of $\phi$. In the two examples above one can take ϕ=italic-$\phi=$ (position, momentum) and ϕ=italic-$\phi=$ full spin vector. In the last example, say in the spin 1/2 case, one can model the discrete spin component $\theta^{a}$ in direction $a$ as $\mathrm{sign}(\mathrm{cos}(a,\phi))$. Giving $\phi$ a reasonable distribution here, results in a correct distribution of each $\theta^{a}$.
@@
In this sense, one can always think of some relation between $\theta$ and $\eta$ even when they are different maximal accessible variables. Variables that are maximal, different and not in one-to-one correspondence, are called - following Niels Bohr - for complementary.
@@
The second step is to show the following: If $k$ is the transformation connecting two related variables $\theta$ and $\eta$, then there is a unitary matrix $W(k)$ such that $A^{\eta}=W(k)^{-1}A^{\theta}W(k)$. (A matrix $W$ is unitary if $W^{-1}=W^{\dagger}$, where the last matrix is obtained from $W$ by transposing it and taking the complex conjugate of all its elements.)
Given these results, a rich theory follows. The set of eigenvalues of the operator $A^{\theta}$ is identical to the set of possible values of $\theta$. The variable $\theta$ is maximal if and only if all eigenvalues of the corresponding operator are simple. In general, the eigenspaces of $A^{\theta}$ are in one-to-one correspondence with questions ‘What is $\theta$’/ ‘What will $\theta$ be if we measure it?’ together with sharp answers $\theta=u$, where $u$ is an eigenvalue of $A^{\theta}$.
@@
A Hilbert space $\mathcal{H}$ can be taken as an $r$-dimensional complex vector space. Operators on this space are then $r\times r$ matrices. An operator $A$ is said to be self-adjoint if $A^{\dagger}=A$, where $A^{\dagger}$ is formed from $A$ by transposing it and taking complex conjugates of all elements. Self-adjoint operators have real-valued eigenvalues.
A ket vector $|\psi\rangle$, a state vector, can be seen as a unit column vector in $\mathcal{H}$. I will concentrate mainly on state vectors of the form $|\psi\rangle=|b;j\rangle$, an eigenvector of an operator $A^{b}$, in my language, equivalent to a particular answer to a question of the form ‘What is $\theta^{b}$?’, where $\theta^{b}$ is a maximal accessible theoretical variable.
@@
The projection operator $\Pi_{j}^{b}=|b;j\rangle\langle b;j|$ projects onto the state vector $|b;j\rangle$. In general, a projection operator $\Pi_{B}$ projects onto the plane $B$.
@@ 2.2 A simple model of a person’s decisions
Consider a person $C$ in some decision situation. Say that he or she has the choice between a finite set of actions $a_{1},...,a_{r}$. Relative to this situation we can define a finite-valued decision variable $\theta$, taking the different values $1,...,r$, such that $\theta=i$ corresponds to the action $a_{i}$ $(i=1,...,r)$. If $C$ really is able to make a decision here and carry out the actions, we say that $\theta$ is an accessible variable. In analogy with the situation in the previous subsection, the variable $\theta$ is in relation to a person $C$, in fact, here $\theta$ belongs to the mind of $C$.
@@
This must be made precise. A decision problem is said to be maximal if $C$ is just able to make his mind with respect to this decision; if the problem is made slightly more complicated, he is not able to take a decision. Let two different maximal decision variables be $\theta$ and $\eta$, where $\theta=i$ corresponds to the action $a_{i}$ $(i=1,...,r)$, and $\eta=j$ corresponds to the action $b_{j}$ $(j=1,...,r)$. Then, by the theory of the previous subsection, we can model the situation by using quantum theory.
@@
Another simple model is the following: Let $\eta$ be a maximal accessible of some kind, describing the relevant state of the observer $C$, and let $\theta$ be a maximal accessible decision variable. Theorem 1 is equally valid in this situation.
In this article, I will mainly focus on maximal decision problems. A thorough discussion including decisions that are not maximal, is given in Mogiliansky et al. (2009).
2.3 When is a macroscipic accessible variable maximal?
For the purpose of this article, it is crucial that my foundation of quantum theory also is valid when relevant accessible variables are macroscopic. In the language of Khennikov (2010), I also want to include quantum-like models, which are applicable in biology, economy, psychology and many other disciplines. Quantum structure may be ubiquitous. A special case is quantum cognition theory (Wang et al., 2013; Pothos and Busemeyer, 2022), which includes quantum decision theory.
The crucial assumption of Theorem 1 is that there exist two different maximal accessible theoretical variables in some given situation. This situation may well be macroscopic.
I will start with an example from Helland (2021). In a medical experiment, let $\mu_{a},\mu_{b},\mu_{c}$ and $\mu_{d}$ be continuous inaccessible parameters, the hypothetical effects of treatment $a,b,c$ and $d$, respectively. Assume that the focus of the experiment is to compare the effect of treatment $b$ with the mean effect of the other treatments, which is supposed to give the parameter $\frac{1}{3}(\mu_{a}+\mu_{c}+\mu_{d})$. One wants to do a pairwise experiment, but it turns out that the maximal parameter that can be estimated is $\theta^{b}=\mathrm{sign}(\mu_{b}-\frac{1}{3}(\mu_{a}+\mu_{c}+\mu_{d}))$.
In the language of the present article, $\theta^{b}$ here is a maximal accessiblble theoretical variable. In concrete terms, one can imagine the following situation: One has four different ointments against rash. A patient is treated with ointment $b$ on one side of his back, a mixture of the other ointments on the other side of his back. It is only possible to observe which side improves best but this observation is assumed to be very accurate. In practice the experiment is done on many patients, and one selects out the patients were the difference is clear.
The concrete factors above make $\theta^{b}$ maximal. For a complete description of the experiment, see Helland (2021), Subsection 5.6.4. The application of a quantum-like model to the important field of experimental design is a completely new area that ought to be treated more thoroughly.
Several examples are given in Haven and Khrennikov (2013), say, and similar discussions could have been made with some of these examples. I will maintain, however, that the term maximal has a precise definition, as given in Subsection 2.1. Let first a partial ordering among theoretical variables be defined by $\alpha\leq\beta$ if $\alpha=f(\beta)$ for some function $f$. An accessible variable $\theta$ is maximal if it is maximal among the accessible variables with respect to this partial ordering.
In a setting with decision variables, such a variable can be seen to be maximal if the corresponding actor(s) just is/are able to make the decision.
@@ 2.4 The Hilbert space formalism for decisions
In Helland (2022c), Theorem 2 says essentially: Imagine a person $C$ which in some context has two related maximal accessible variables $\theta$ and $\eta$ in his mind. Impose a specific symmetry assumption. Then $C$ cannot simultaneous have in mind any other maximal accessible varable which is related to $\theta$, but not related to $\eta$. It was claimed in Helland (2022c) that the violation of a famous inequality by practical Bell experiments, can be understood on the basis of this theorem. See also Helland (2023b), where a corresponding theorem is formulated without any symmetry assumption.
Note that this result has the qualification ‘at the same time’, and indicate a specific restriction to two maximal variables. But the human mind is very flexible. Taking time into account, we can think of very many variables, even ones that are not related.
@@ 2.5 The Born rule; general arguments
Born’s formula is the basis for all probability calculations in quantum mechanics. In textbooks it is usually stated as a separate axiom, but it has also been argued for by using various sets of assumptions; see Campanella et al. (2020) for some references. In fact, the first argument for the Born formula, assuming that there is an affine mapping from set of density functions to the corresponding probability functions, is due to von Neumann (1927). Here we will try to use assumptions that are as weak as possible, and assumptions that can be related to notions both from statistical theory and quantum theory. We start with assuming the likelihood principle: The experimental evidence from any experiment must be based upon the likelihood $l$, which is probability density or the point probability of observations, seen as a function of the full parameter.
@@
In Helland (2021), a generalized likelihood principle is proved from the ordinary likelihood principle: Given some experiement, or more generally, some context $\tau$ connected to an experiment, any experimental evidence will under weak assumptions be a function of the likelihood effect $F$. In particular, the relevant probabilities $q$ sre functions of $F$: $q(F|\tau)$.
@@
[…]
@@ 2.6 A survey of quantum probabilities
$=\|\Pi_{i}^{c}\frac{\Pi_{j}^{b}|\psi\rangle}{\|\Pi_{j}^{b}|\psi\rangle\|}\|^{2}\|\Pi_{j}^{b}|\psi\rangle\|^{2}=\|\Pi_{i}^{c}\Pi_{j}^{b}|\psi\rangle\|^{2}.$
@@ 2.7 Born’s rule in a decision setting
Go back to the person $C$ who is facing at least two related maximal decisions, with corresponding decision variables $\theta$ and $\eta$. Each of these decisions may be composed of several partial decisions; then the decision variables may be thought of as vectors. However, we will make the simplifying assumption that each partial decision has a finite number of outcomes, say possible actions; then the total decision also has a finite number of outcomes, and without loss of generality, we can order these actions as $\{a_{i};i=1,2,..,r\}$.

### v5

@@ Abstract
Quantum decision theory is introduced here, and new basis for this theory is proposed. It is first based upon the author’s general arguments for the Hilbert space formalism in quantum theory, next on arguments for the Born rule, that is, the basis for calculating quantum probabilities. A basic notion behind the quantum theory foundation is that of theoretical variables, that are divided into accessible and inaccessible ones. This is here specialized to decision variables. It is assumed that all accessible variables can be seen as functions of a specific inaccessible variable. Another assumption is that there exist two different maximal accessible theoretical variables in the given situation. Two basic assumptions behind the Born rule are 1) the likelihood principle, 2) the actor in question has motivations that can be modeled by a hypothetical perfectly rational higher being. The theory is illustrated by a medical example. Finally, a brief discussion of decision processes is given.
@@ 1 Introduction
In standard economic theory and statistical theory, the problem of optimizing behaviour is reduced to optimizing expected reward or utility. I will claim that this does not quite correspond to the way most people act during a decision process. This statement can be supported by a number of empirical investigations; see for instance the review articles mentioned below in connection to quantum decision theory. The argument will be further elucidated later in this article.
@@
A rather primitive form of learning is reinforcement learning, learning governed by rewards or punishments. In Friston et al. (2009) the notion of active inference was introduced as an alternative to reinforcement learning in connection to decision making. This may be important, since the way many people learn in connection to decisions obviously is more active than what can be described as reinforcement learning. In op cit. a theory of active inference is formalized in terms of a free-energy formulation of perception, quantizised in terms of entropy or average surprise associated with a probability distribution on an agent’s state space and his environment.
@@
Theories of decisions can take many points of departure. In this article we will describe a new foundation of a comparatively recent theory: Quantum decision theory (QDT). This foundation is completely based upon a new foundation of quantum mechanics, discussed in the book Helland (2021) and in some newer articles, Helland (2022a,b,c, 2023a,b,c, 2024a,b,c, 2025a,b, 2026).
@@
The plan of this article is as follows: I first review briefly the new approach to the foundation of quantum theory, and then I use this to derive and motivate a new foundation of quantum decision theory, taking as a basis a simple model of a person’s decisions. First, the Hilbert space formalism of decision variables is derived. Then the Born formula, the basis for computing quantum probabilities, is argued for under weak assumptions, and this is specialized to a decision setting. An example of applications is then given, and used to illustrate some of the peculiarities of quantum probabilities. Then my foundation of QDT is further discussed from several points of view before I give some concluding remarks.
@@ 2.1 On the foundation of quantum theory in general
A completely new approach towards quantum foundations is proposed in Helland (2021, 2022a,b,c 2023a,b,c,2024a,b,c, 2025a). The basis of this approach is an observer who is in some physical situation. In this situation there are theoretical variables, and certain of these variables, say $\theta,\lambda,\eta,$... are related to the observer $C$. Some of the variables are accessible to him, meaning roughly that it is, in some future, in principle possible to obtain as accurate values as he wishes on the relevant variable. Others are inaccessible. Examples of the latter are the vector (position, momentum) of a particle at some time, or the full spin vector of a spin particle, an imagined vector whose discretized projection in the direction $a$ is the spin component in that direction. The terms ‘accessible’ and ‘inaccessible’ are primitive notions of the theory.
The main assumption of my theory is then as follows: Related to an observer $C$ there is an inaccessible variable $\phi$ such that all accessible variables can be seen as functions of $\phi$. In the two examples above one can take $\phi=$ (position, momentum) and $\phi=$ full spin vector. In the last example, say in the spin 1/2 case, one can model the discrete spin component $\theta^{a}$ in direction $a$ as $\mathrm{sign}(\mathrm{cos}(a,\phi))$. Giving $\phi$ a reasonable distribution here, results in a correct distribution of each $\theta^{a}$. A symmetry assumption, needed for continuous variables, is that there exists a group $K$ acting on the range of $\phi$.
@@
In this sense, one can always think of some relation between $\theta$ and $\eta$ even when they are different maximal accessible variables. Variables that are maximal, different, have similar range spaces, and not in one-to-one correspondence, are called - following Niels Bohr - for complementary.
@@
The second step is to show the following: If $k$ is the transformation connecting two related variables $\theta$ and $\eta$, then there is a unitary matrix $W(k)$ such that $A^{\eta}=W(k)^{-1}A^{\theta}W(k)$.
Given these results, a rich theory follows. The set of eigenvalues of the operator $A^{\theta}$ is identical to the set of possible values of $\theta$. The variable $\theta$ is maximal if and only if all eigenvalues of the corresponding operator are simple. In general, the eigenspaces of $A^{\theta}$ are in one-to-one correspondence with questions ‘What is $\theta$’/ ‘What will $\theta$ be if we measure it?’ together with sharp answers $\theta=u$, where $u$ is the relevant eigenvalue of $A^{\theta}$.
@@
A Hilbert space $\mathcal{H}$ can be taken as an $r$-dimensional complex vector space. Operators on this space are then $r\times r$ matrices. An operator $A$ is said to be self-adjoint if $A^{\dagger}=A$, where $A^{\dagger}$ is formed from $A$ by transposing it and taking complex conjugates of all elements. Self-adjoint operators have real-valued eigenvalues. An operator $U$ is unitary if $U^{-1}=U^{\dagger}$.
A ket vector $|\psi\rangle$, a state vector, can be seen as a unit column vector in $\mathcal{H}$. I will concentrate on state vectors of the form $|\psi\rangle=|b;j\rangle$, an eigenvector of some operator $A^{b}$, in my language, equivalent to a particular answer to a question of the form ‘What is $\theta^{b}$?’, where $\theta^{b}$ is a maximal accessible theoretical variable.
@@
The projection operator $\Pi_{j}^{b}=|b;j\rangle\langle b;j|$ projects onto the state vector $|b;j\rangle$. In general, a projection operator $\Pi_{B}$ projects onto the line, plane or hyperplane $B$.
@@ 2.2 A simple model of a person’s decisions
Consider a person $C$ in some decision situation. Say that he or she has the choice between a finite set of actions $a_{1},...,a_{r}$. Relative to this situation we can define a finite-valued decision variable $\theta$, taking the different values 1,,r1,...,r, such that $\theta=i$ corresponds to the action $a_{i}$ (i=1,,r)(i=1,...,r). If $C$ really is able to make a decision here and carry out the actions, we say that $\theta$ is an accessible variable. In analogy with the situation in the previous subsection, the variable $\theta$ is in relation to a person $C$, in fact, here $\theta$ belongs to the mind of $C$.
I will regard accessible decision variables as special accessible variables. They may be functions of other theoretical variables that are not necessarily decision variables. But note that any decision variable, $\theta$ say, corresponding to possible actions $a_{1},...,a_{r}$, may be seen as a function of another decision variable: Just expand the set of possible actions. And there are other decision variables that are functions of $\theta$: Collect the actions $\{a_{i}\}$ into subsets of these actions.
In the general theory of Subsection 2.1, every accessible variable $\theta$ is assumed to be a function of a maximal accessible variable $\eta$. When $\theta$ is a decision variable, we can, without loss of generalty assume that this $\eta$ also is a decision variable.
Lemma 1 Let $\theta$ be an accessible decision variable. Then there exists a maximal accessible decision variable $\eta$ and a decision variable $\zeta$ which is a one-to-one function of $\theta$ such that $\zeta\leq\eta$ in the partial ordering defined by functional dependence.
Proof. By assumption, there exists a maximal variable $\xi$ such that $\theta\leq\xi$, that is, $\theta$ is a function of $\xi$. The crucial point is that $\xi$ always can be seen as a one-to-one function of a decision variable $\eta$: Without loss of generality, let the different values of $\xi$ be 1,,s1,...,s for some $s\geq r$, the different actions in $\theta$. Define a decision process for $C$ by taking some arbitrary action $b_{j}$ when $\xi=j$, (j=1,,s)(j=1,...,s). This defines the decision variable $\eta$. Since we must have $s\geq r$, one can collect the actions $b_{j}$ in subsets that are in one-to-one correspondence with $\{a_{i};i=1,...,r\}$. This defines the decision variable $\zeta$.
@@
This must be made precise. A decision problem is said to be maximal if $C$ is just able to make his mind with respect to this decision; if the problem is made slightly more complicated, he is not able to take a decision. Let two different maximal decision variables be $\theta$ and $\eta$, where $\theta=i$ corresponds to the action $a_{i}$ (i=1,,r)(i=1,...,r), and $\eta=j$ corresponds to the action $b_{j}$ (j=1,,r)(j=1,...,r). Then, by the abstract mathematical version of the theory of the previous Subsection, we can model the situation by using quantum mechanics.
@@
Another simple model is the following: Let $\eta$ be a maximal accessible variable of some kind, describing the relevant state of the observer $C$. For simplicity, assume that $\eta$ takes a finite number $s$ of values. Let $\theta$ be a maximal accessible decision variable taking $r$ values, which can be extended to $s$ by adding some impossible actions. Theorem 1 is equally valid in this situation.
In this article, I will mainly focus on maximal decision problems. A thorough discussion including decisions that are not maximal, is given in Mogiliansky et al. (2009). By Lemma 1, each decision variable may be seen as a function of a maximal accessible decision variable.
I will maintain that the term maximal has a precise definition, as given in Subsection 2.1. To repeat , let first a partial ordering among theoretical variables be defined by $\alpha\leq\beta$ if $\alpha=f(\beta)$ for some function $f$. An accessible variable $\theta$ is maximal if it is maximal among the accessible variables with respect to this partial ordering.
In a setting with decision variables, such a variable can be seen to be maximal if the corresponding actor(s) just is/are able to make the decision. If one adds one extra possible action, $C$ is then unable to make the decision.
@@ 2.3 The Hilbert space formalism for decisions
In Helland (2022c), Theorem 2 says essentially: Imagine a person $C$ which in some context has two related maximal accessible variables $\theta$ and $\eta$ in his mind. Impose a specific symmetry assumption. Then $C$ cannot simultaneous have in mind any other maximal accessible varable which is related to $\theta$, but not related to $\eta$. It was claimed in Helland (2022c) that the violation of a famous inequality by practical Bell experiments, can be understood on the basis of this theorem. See also Helland (2023b), where a corresponding theorem is formulated without any symmetry assumption in the case where the maximal accessible variables take a finite number of values.
Note that this result has the qualification ‘at the same time’, and indicate a specific restriction to the two maximal variables. But the human mind is very flexible. Taking time into account, we can think of very many variables, even ones that are not related.
@@ 2.4 The Born rule; general arguments
Born’s formula is the basis for all probability calculations in quantum mechanics. In textbooks it is usually stated as a separate axiom, but it has also been argued for by using various sets of assumptions; see Campanella et al. (2020) for some references. In fact, the first argument for the Born formula, assuming that there is an affine mapping from set of density functions to the corresponding probability functions, is due to von Neumann (1927). Here we will try to use assumptions that are as weak as possible, and assumptions that can be related to notions both from statistical theory and quantum theory. We start with assuming the likelihood principle: The experimental evidence from any experiment must always be based upon the likelihood $l$, which is probability density or the point probability of observations, seen as a function of the full parameter.
@@
In Helland (2021), a generalized likelihood principle is proved from the ordinary likelihood principle: Given some experiement, or more generally, some context $\tau$ connected to an experiment, any experimental evidence will under weak assumptions be a function of the likelihood effect $F$. In particular, the relevant probabilities $q$ are functions of $F$: $q(F|\tau)$.
The following Theorem and its Corollary are proved in Helland (2021):
@@
$\begin{array}[]{l}(1)\ 0\leq\mu(F)\leq 1\ \mathrm{for\ all}\ F,\\ (2)\ \mu(I)=1,\\ (3)\ \mu(F_{1}+F_{2}+\ldots)=\mu(F_{1})+\mu(F_{2})+\ldots\ \mathrm{whenever}\ F_{1}+F_{2}+\ldots\leq I.\end{array}$
@@ 2.5 A survey of quantum probabilities
$\|\Pi_{i}^{c}\frac{\Pi_{j}^{b}|\psi\rangle}{\|\Pi_{j}^{b}|\psi\rangle\|}\|^{2}\|\Pi_{j}^{b}|\psi\rangle\|^{2}=\|\Pi_{i}^{c}\Pi_{j}^{b}|\psi\rangle\|^{2}.$
@@ 2.6 Born’s rule in a decision setting
Go back to the person $C$ who is facing at least two different maximal decisions, with corresponding decision variables $\theta$ and $\eta$. Each of these decisions may be composed of several partial decisions; then the decision variables may be thought of as vectors. However, we will make the simplifying assumption that each partial decision has a finite number of outcomes, say possible actions; then the total decision also has a finite number of outcomes, and without loss of generality, we can order these actions as $\{a_{i};i=1,2,..,r\}$.